OpenStudy (anonymous):

A kilobyte is 2 to the power of 10. A megabyte is 2 to the power of 20 bytes. How many kilobytes are in a megabyte? Use the exponent rule for division to find your answer and give your answer in exponential form.

OpenStudy (anonymous):

\[x^a * x^b = x^{a+b}\] \[2^{10} * X = 2^{20}\]

OpenStudy (anonymous):

what is X????????

OpenStudy (anonymous):

the number of kilobytes in a megabyte. X = 2^10 since \[2^{10}*2^{10} = 2^{20}\]

OpenStudy (anonymous):

its like saying i spent 8$ on apples that were 2$ each. how many did i buy? 2 * X = 8

OpenStudy (anonymous):

its like saying i spent 8$ on apples that were 2$ each. how many did i buy? the equation is: 2 * X = 8. and one would solve for X. X = 4 obv